# An FPGA-Based Open-Source Hardware-Software Framework for Side-Channel Security Research

### v1

@@ III-B3 Host PC
The host PC drives the whole proposed framework through the flow software scripts, which manage the interaction with both hardware devices, namely the prototype board and the oscilloscope, and software tools, such as the EDA and compilation toolchains and state-of-the-art SCA attack scripts. Matching the high-level flow previously described in Section III-A, we can identify three main top-level scripts that drive the corresponding phases of the framework.
@@ IV Microarchitecture
IV-A Debug subsystem
IV-A1 Global and local debug units
Two local debug units are instantiated for the CPU core and for the DFS actuator, respectively, to act as adapters between the global debug unit and the two bus masters. The local debug unit interface with the global one is common to all the local debug units, while the interface with the CPU or DFS actuator is custom tailored to the specific interactions that are implemented with such module. The CPU local debug unit can access the program counter (PC), registers, and performance monitoring counters (PMCs) of the CPU, as well as halting, resetting, and restarting it and advancing its execution by a single step. The DFS one can set a new target clock frequency for the DFS actuator, get the frequency of the current clock signal generated and configure the DFS to randomly switch the frequency of the clock signal. Figure 3 depicts the microarchitecture of the debug subsystem of the SoC, highlighting the global debug unit, the CPU and DFS local debug units, and their interactions with each other as well as with the CPU and the DFS actuator.
@@ IV-C DFS actuator
The DFS actuator, whose microarchitecture is depicted in Figure 5, leverages two mixed-mode clock manager (MMCM) components to provide a glitch- and latency-free switching between different clock frequencies at run time. It receives as its inputs a target clock frequency (fin in Figure 5), a 1-bit flag to enable random DFS (rnd), and a reference clock signal (clkref), and it outputs the generated clock signal (clkout) and the current clock frequency (fout). The rnd flag selects the actual target clock frequency for the DFS reconfiguration between the one received through the fin input by the DFS local debug unit and a random value frnd output by an internal TRNG with similar to the one in Section IV-B. Such target clock frequency, registered to be output to the DFS local debug unit and readable through a dedicated debug command, is decoded and used as an address to select a corresponding set of MMCM configuration parameters from the BRAM memory.
@@ V Experimental Evaluation
The framework is released open source, including all the hardware and software that comprise it. The open-source nature of the framework, as well as the adoption of standard languages ranging from SystemVerilog for the RTL description of the SoC and the testbench to C for the application and RTOS sources and Python for the flow scripts, enable the user to expand its capabilities and support novel SCA attacks and countermeasures. The goal of the proposed framework is not introducing new SCA attack or countermeasure techniques, but providing instead an infrastructure that implements the current state of the art and thus enables new research in the field. The experimental campaign described in this section aims to showcase how the proposed framework delivers an effective research platform for SCA attacks and countermeasures on FPGA targets. Section V-A lists the hardware and software required by the framework, Section V-B details the specific setup employed for the experimental evaluation, Section V-C delivers an experimental analysis of the SCA techniques provided out of the box.
@@ V-A Hardware and software requirements
The framework makes use of widely available devices and tools, including the development board and FPGA chips for prototyping, the digital oscilloscope for power trace measurement, and the software required for the various phases of the flow ranging from EDA synthesis and place-and-route to compilation and SCA resistance evaluation.
@@ V-A1 Prototype
XDC design constraint files are provided for prototyping on the NewAE Technology CW305 Artix FPGA Target board, specifically designed for power analysis and fault injection attacks against hardware cryptographic functions [37]. The CW305 board mounts a BGA socket that can accommodate any chip from the AMD Artix-7 mid-range FPGA family, which is the most widely adopted in academic and industrial research, including the security and cryptography fields [38]. The proposed SoC fits into AMD Artix-7 50 chips, which pack 32600 look-up tables (LUT), 65200 flip-flops (FF), 120 digital signal processing (DSP) blocks, and 75 36kb blocks of block RAM (BRAM), and larger ones from the same family. Simulation of the virtual prototype is supported by the AMD Vivado EDA toolchain.
@@ V-A2 Oscilloscope
Power measurements are carried out through an oscilloscope from the Pico Technology PicoScope 5000 Series family, which features a 200MHz maximum bandwidth and can collect up to 1 billion samples per second at a resolution ranging from 8 to 16 bits [39]. The oscilloscope has its A analog channel connected to the board through a coaxial SMA connector to measure the FPGA voltage, while its B analog channel is connected to a general-purpose I/O pin that outputs the trigger signal which marks the start and the end of the computation.
@@ V-B3 Breakpoints and triggerpoints
V-C Experimental analysis
The experimental analysis applies to the computing platform configured as described in Section V-B the SCA countermeasure and attack techniques from the state of the art included in the framework, quantitatively assessing their effectiveness.
V-C1 SCA countermeasure techniques
Clock frequency randomization
@@ V-C1 SCA countermeasure techniques
The morphing countermeasure leverages the TRNG to randomly modify the execution of a cryptographic operation [31]. In the use case considered in the experimental evaluation, morphing targets the AddRoundKey and SubBytes steps of the AES cryptosystem. Each AddRoundKey execution makes use of a randomly selected version of the XOR operation, chosen among 8 equivalent implementations with different power consumption profiles. SubBytes executions are instead morphed by employing two S-Boxes that are masked by random values and periodically refreshed.
@@
The TRNG component and the support for software multithreading provided by FreeRTOS enable the adoption of the chaff approach, described in [40] and whose pseudocode is listed in Algorithm 1. The latter technique executes, concurrently to a thread for the actual instance of a cipher (line 2 in Algorithm 1), e.g., the AES encryption of a plaintext with a certain key, a set of additional threads performing the same cryptographic operation on the same input but with different keys, properly generated to be correlated with the original one (lines 3–5). All the threads are started and executed in parallel, according to a random scheduler that leverages the TRNG instantiated in the SoC (line 6), until the thread performing the actual cryptographic operation has completed (lines 7–10).
@@ V-C2 SCA attack techniques
The SCA attacks provided out of the box by the framework include 1) a CPA attack [7], that extracts the secret key of the target cryptosystem executions by correlating their power signature with their operating behavior, performed in the experimental campaign on up to 1 million AES power traces, 2) a template profiled attack [9], that leverages the Bayes’ theorem to estimate the probability of a key given its multivariate normal distribution, carried out by using 1024 traces for each of the 256 possible AES key bytes during both the profiling and attack phases, and 3) an ML-based attack leveraging convolutional neural networks (CNNs) [32], that learn a recurrent pattern of the leakage to predict the secret key and whose training and inference are carried out on the same dataset as the template attack.
@@ V-C3 SCA security assessment
The experimental analysis assesses the vulnerability of the computing platform to the previously described attacks while executing the AES cryptosystem, whose source code is included in the framework. In particular, we evaluate whether each attack, specifically targeting the SubBytes step of the first AES round, succeeds or fails when executing the plain AES cryptosystem without any protection, a software-masked version of AES, and plain AES protected by the clock frequency randomization, morphing, and chaff SCA hiding countermeasures.
TABLE II: Quality metrics obtained by attacks on the experimental platform when implementing different countermeasures. Legend: – undefined.
@@
More in detail, clock frequency randomization and chaff are the most effective countermeasures against the proposed SCA attacks, as they achieve a high guessing entropy and a success rate close to 0% in every scenario. The morphing countermeasure is instead shown to be more vulnerable, as every technique can break it. The CNN attack can isolate the correct key with a guessing distance of 0.81, using only eight traces, while CPA and template attacks are less effective though successful. Guessing distance values are notably undefined for the CPA attack, since it is not a profiled attack.
VI Conclusions
This manuscript introduced a novel open-source framework for research on SCA targeting FPGA-based IoT/microcontroller-class computing platforms. The framework includes a RISC-V-based IoT-class SoC that features an ad-hoc debug infrastructure to maximize the observability and controllability of the computing platform and thus simplify the execution of SCA attacks, as well as a DFS actuator, a TRNG, and a timer that provide support for a set of state-of-the-art SCA countermeasures available out of the box. A complete automated flow encompasses the configuration of the SoC, the execution of target applications and corresponding collection of side-channel information, and the analysis to identify eventual SCA vulnerabilities and pinpoint the sources of side-channel information leakage.
@@ VI Conclusions
Future developments foresee the addition of a multi-core CPU architecture and cryptographic hardware accelerators and support for fault attacks and countermeasures against them.

### v2

@@ III-B3 Host PC
The host PC drives the whole JARVIS framework through the flow software scripts, which manage the interaction with both hardware devices, namely the prototype board and the oscilloscope, and software tools, such as the EDA and compilation toolchains and state-of-the-art SCA attack scripts. Matching the high-level flow previously described in Section III-A, we can identify three main top-level scripts that drive the corresponding phases of the framework.
@@ IV Microarchitecture
IV-A Debug subsystem
IV-A1 Global and local debug units
Two local debug units are instantiated for the CPU core and for the DFS actuator, respectively, to act as adapters between the global debug unit and the two bus masters. The local debug unit interface with the global one is common to all the local debug units, while the interface with the CPU or DFS actuator is custom tailored to the specific interactions that are implemented with such module. The CPU local debug unit can access the program counter (PC), registers, and performance monitoring counters (PMCs) of the CPU, as well as halt, reset, and restart it and advance its execution by a single step. The DFS one can set a new target clock frequency for the DFS actuator, get the frequency of the current clock signal generated and configure the DFS to randomly switch the frequency of the clock signal. Figure 3 depicts the microarchitecture of the debug subsystem of the SoC, highlighting the global debug unit, the CPU and DFS local debug units, and their interactions with each other as well as with the CPU and the DFS actuator.
@@ IV-C DFS actuator
The DFS actuator, whose microarchitecture is depicted in Figure 5, leverages two mixed-mode clock manager (MMCM) components to provide a glitch- and latency-free switching between different clock frequencies at run time [17]. It receives as its inputs a target clock frequency (fin in Figure 5), a 1-bit flag to enable random DFS (rnd), and a reference clock signal (clkref), and it outputs the generated clock signal (clkout) and the current clock frequency (fout). The rnd flag selects the actual target clock frequency for the DFS reconfiguration between the one received through the fin input by the DFS local debug unit and a random value frnd output by an internal TRNG with similar to the one in Section IV-B. Such target clock frequency, registered to be output to the DFS local debug unit and readable through a dedicated debug command, is decoded and used as an address to select a corresponding set of MMCM configuration parameters from the BRAM memory.
@@ V Experimental Evaluation
The experimental campaign aims to showcase how the JARVIS framework delivers an effective research platform for SCA attacks and countermeasures on FPGA targets. Section V-A lists the hardware and software required by the framework, Section V-B details the specific setup employed for the experimental evaluation, and Section V-C delivers an analysis of the SCA techniques provided out of the box.
@@ V-A Hardware and software requirements
The framework makes use of widely available devices and tools, including the development board and FPGA chips for prototyping, the digital oscilloscope for power trace measurement, and the software required for the various phases of the flow, ranging from EDA synthesis and place-and-route to compilation and SCA resistance evaluation.
@@ V-A1 Prototype
XDC design constraint files are provided for prototyping on the NewAE Technology CW305 Artix FPGA Target board, specifically designed for power analysis and fault injection attacks against hardware cryptographic functions [^2] More information on NewAE Technology CW305 Artix FPGA Target available at https://rtfm.newae.com/Targets/CW305%20Artix%20FPGA.. The CW305 board mounts a BGA socket that can accommodate any chip from the AMD Artix-7 mid-range FPGA family, which is the most widely adopted in academic and industrial research, including the security and cryptography fields [43]. The proposed SoC fits into AMD Artix-7 50 chips, which pack 32600 look-up tables (LUT), 65200 flip-flops (FF), 120 digital signal processing (DSP) blocks, and 75 36kb blocks of block RAM (BRAM), and larger ones from the same family.
@@ V-A2 Oscilloscope
Power measurements are carried out through an oscilloscope from the Pico Technology PicoScope 5000 Series family, which features a 200MHz maximum bandwidth and can collect up to 1 billion samples per second at a resolution ranging from 8 to 16 bits [^3] More information on Pico Technology PicoScope 5000 Series available at https://www.picotech.com/oscilloscope/5000/picoscope-5000-specifications.. The oscilloscope has its A analog channel connected to the board through a coaxial SMA connector to measure the FPGA voltage, while its B analog channel is connected to a general-purpose I/O pin that outputs the trigger signal which marks the start and the end of the computation.
@@ V-B3 Breakpoints and triggerpoints
V-C Experimental analysis
The experimental analysis applies to the computing platform, configured as described in Section V-B, the SCA countermeasure and attack techniques from the state of the art included in the framework, quantitatively assessing their effectiveness.
V-C1 SCA countermeasure techniques
Clock frequency randomization
@@ V-C1 SCA countermeasure techniques
The morphing countermeasure leverages the TRNG to randomly modify the execution of a cryptographic operation [36]. For example, in the AES use case considered in the experimental evaluation, morphing targets the AddRoundKey and SubBytes steps of the AES cryptosystem. Each AddRoundKey execution employs a randomly selected version of the XOR operation, chosen among 8 equivalent implementations with different power consumption profiles. SubBytes executions are instead morphed by using S-Boxes that are masked by random values and periodically refreshed.
@@
The TRNG component and the support for software multithreading provided by FreeRTOS enable the adoption of the chaff approach, described in [44] and whose pseudocode is listed in Algorithm 1. The latter technique executes, concurrently to a thread for the actual instance of a cipher (line 2 in Algorithm 1), e.g., the encryption of a plaintext with a certain key, a set of additional threads performing the same cryptographic operation on the same input but with different keys, properly generated to be correlated with the original one (lines 3–5). All the threads are started and executed in parallel, according to a random scheduler that leverages the SoC’s TRNG (line 6), until the thread performing the actual cryptographic operation has completed (lines 7–10).
@@ V-C2 SCA attack techniques
The SCA attacks provided out of the box by the framework include 1) a CPA attack [7], that extracts the secret key of the target cryptosystem executions by correlating their power signature with their operating behavior, performed in the experimental campaign on up to 1 million power traces, 2) a template profiled attack [9], that leverages the Bayes’ theorem to estimate the probability of a key given its multivariate normal distribution, carried out by using 1024 traces for each of the 256 possible key bytes during both the profiling and attack phases, and 3) an ML-based attack leveraging convolutional neural networks (CNNs) [37], that learn a recurrent pattern of the leakage to predict the secret key and whose training and inference are carried out on the same dataset as the template attack.
@@ V-C3 SCA security assessment
The experimental analysis assesses first the vulnerability of the computing platform to the previously described attacks while executing the AES-128 cryptosystem. We make use of the constant-time, S-box-based AES-128 C implementation from the widely adopted OpenSSL [45] cryptography suite. In particular, we evaluate whether each attack, specifically targeting the SubBytes step of the first AES round, succeeds or fails when executing the plain AES cryptosystem without any protection, a software-masked version of AES, and plain AES protected by the clock frequency randomization, morphing, and chaff SCA hiding countermeasures.
TABLE II: Quality metrics obtained by attacks to AES execution on the experimental platform when implementing different countermeasures. Undefined values are denoted by –.
@@
More in detail, clock frequency randomization and chaff are the most effective countermeasures against the proposed SCA attacks, as they achieve a high guessing entropy and a success rate close to 0% in every scenario. The morphing countermeasure is instead shown to be more vulnerable, as every technique can break it. The CNN attack can isolate the correct key with a guessing distance of 0.81, using only eight traces, while CPA and template attacks are less effective though successful. Guessing distance values are notably undefined for the CPA attack, since the latter is not a probabilistic attack. Guessing distance measures indeed a probability distance and, thus, it is suitable only for attacks such as template and CNN ones that compute the probability of a key guess being correct.
@@
TABLE III: Quality metrics obtained by template attacks to Camellia, Clefia, and Seed execution when implementing different countermeasures. Undefined values are denoted by –.
Cryptosystem
Countermeasure Quality metric Camellia Clefia Seed
None Guessing entropy 1 1 1
Guessing distance 0.28 0.61 0.23
Success rate 100% 100% 100%
Number of traces 1202 52 1021
Clock frequency Guessing entropy 133.09 105.53 132.63
randomization Guessing distance -0.49 -0.45 -0.50
Success rate 0% 0% 0%
Number of traces
Morphing Guessing entropy 9 1 55.75
Guessing distance -0.06 0.31 -0.31
Success rate 40% 100% 3%
Number of traces 1379
Chaff Guessing entropy 148.06 119.72 133.84
Guessing distance -0.56 -0.48 -0.56
Success rate 0% 0% 0%
Number of traces
To further demonstrate the flexibility of the proposed hardware-software framework, Table III reports the results of the template attack when the computing platform executes three additional cryptographic applications from the OpenSSL [45] suite, namely, the Camellia, Clefia, and Seed block ciphers. For each evaluated cryptosystem, results are reported considering no countermeasures, the use of clock frequency randomization, the use of morphing and the use of chaffing. The reported quality metrics to assess the side-channel vulnerability refer to the SubBytes operation in the first round of each cryptosystem.
The results show that applying morphing to protect Clefia is ineffective against template attacks, whereas it is moderately and highly effective in the Camellia and Seed use cases, respectively. Conversely, the clock frequency randomization and chaff countermeasures are shown to successfully thwart the template attack when applied during the execution of all three cryptographic applications.
VI Conclusions
This manuscript introduced a novel open-source framework for research on SCA targeting FPGA-based IoT-class computing platforms. The framework includes a RISC-V-based IoT-class SoC that features an ad-hoc debug infrastructure to maximize the observability and controllability of the computing platform and thus simplify the execution of SCA attacks, as well as a DFS actuator, a TRNG, and a timer that provide support for a set of state-of-the-art SCA countermeasures available out of the box. A complete automated flow encompasses the configuration of the SoC, the execution of target applications and corresponding collection of side-channel information, and the analysis to identify eventual SCA vulnerabilities and pinpoint the sources of side-channel information leakage.
@@ VI Conclusions
Future developments foresee the addition of a dual-core CPU architecture and of accelerators specifically dedicated to cryptography purposes. Such optional features, limited in terms of additional resources to still deliver a lightweight IoT-class computing platform, will not compromise the observability and controllability of the system for SCA, which are essential for detecting side-channel leakage and its sources. In addition, we plan to extend our framework to support research on fault attacks and countermeasures against the latter.